(i.) What is meant by Doppler's effect?

  (ii.) A bat flies at a steady speed of 5ms-1 and emits waves of frequency 85.0kHz. What is the

       frequency of the wave detected by the bat after reflection from a stationary obstacle directly ahead of the bat?


1
Expert's answer
2021-03-09T15:31:14-0500

i) Doppler Effect refers to the change in wave frequency during the relative motion between a wave source and its observer. For instance, when a sound object moves towards you, the frequency of the sound waves increases, leading to a higher pitch.


ii)


"f_o=\frac{c}{c-v_s}f_s=\frac{340}{340-5}85000=86269\ Hz"


"f_o=\left(1+\frac{v}{c}\right)f_s=\left(1+\frac{5}{340}\right)86269=87.5\ kHz"